OpenStudy (anonymous):

If sinx = 1/3 and sec y = 5/4, where x and y lie btwn 0 and pi/2, evaluate sin(x+y).

OpenStudy (anonymous):

sin (A + B) = sin A cos B + cos A sin B If sec y = 5/4, then cos y = 4/5. sin (x + y) = sin x cos y + cos x sin y sin (x + y) = (1/3)(4/5) + (2√2/3)(3/5) sin (x + y) = 4/15 + 2√2/5 4 + 6√2 ------------ = sin(x + y) <==ANSWER . 15

OpenStudy (anonymous):

how did u get 2sqtr 2 / 3 as cos x?. for some reason i got sqrt 8/3.

OpenStudy (anonymous):

??????????

myininaya (myininaya):